OpenStudy (anonymous):

A 13 ft ladder is leaning against a house when its base begins to slide away. By the time the base is 3 ft from the house, the base is moving at the rate of 2 ft/sec. Answer the following: At what speed is the top of the ladder sliding down the wall at that time?

OpenStudy (anonymous):

Keep getting -12/(2sqrt160) but that's wrong..

OpenStudy (anonymous):

Do I differentiate x^2+y^2=13^2 or do I differentiate 3^2+y^2=13^2?

OpenStudy (anonymous):

Refer to the attachment from Mathematica 9.

OpenStudy (irishboy123):

 to start H^2 + W^2 = 13^2 then, after slipping in x (to left) and y (down wall) directions: (H-y)^2 + (W+x)^2 = 13^2 differentiate this 2nd equation wrt time (t): 2(H-y)dy/dt + 2(W+x)dx/dt = 0 dy/dt = -dx/dt {(W+x)/(H-y) } plugging in some of the numbers gives dy/dt = -(2) { 3/ (H-y) } = -6 / (H-y) so what is (H - y)? well from the original equation (H-y)^2 + (W+x)^2 = 13^2 H - y = sqrt ( 13^2 - 3 ^ 2) = sqrt (160) = sqrt (16 x 10) = 4 sqrt(10) thus dy/dt = -6sqrt(10)/4 = etc etc

OpenStudy (irishboy123):

slight correction: -2(H-y)dy/dt + 2(W+x)dx/dt = 0 NB the minus sign at the start ==> then cancels the minus sign at the end. dy/dt is positive.....

OpenStudy (phi):

***Keep getting -12/(2sqrt160) but that's wrong.. *** numerically, that seems to be the correct answer take the derivative with respect to time of x^2 + y^2 = 169 Perhaps they want the answer "rationalized" and written as \[ - \frac{3 \sqrt{10}}{20} \]

OpenStudy (irishboy123):

well my answer, 6/4sqrt(10) = 3/2sqrt(10) = 3sqrt(10)/2(10) = 3sqrt(10)/20

OpenStudy (phi):

also, they ask for the speed, so they may expect a positive value. (the sign will depend on where we put the coordinate axes, but in general, speed is a positive quantity)

OpenStudy (anonymous):

Awesome yep I just had to remove the -.... That's annoying hahah I redid the problem like 3 times and couldnt find my error!

OpenStudy (anonymous):

Thanks everyone